OpenStudy (anonymous):

Factor The Quadratic Equation : -6x+1=-9x^2 .

OpenStudy (anonymous):

Show The Work ? Not Understanding .

OpenStudy (anonymous):

(3x-1)^2

OpenStudy (anonymous):

sorry...itz1/3

OpenStudy (anonymous):

see this one is in (a-b)^2 formula so anwer is 1/3

OpenStudy (mimi_x3):

9x^2 -6x+1= 0 (3x-1)^2=0

OpenStudy (anonymous):

9x^2 - 6x + 1 = 0 x^2 - 6/9x + 1/9 = 0 two numbers that mulitply to get 1/9 and add to get -6/9 (x-1/3)(x-1/3) since 2*(-1/3) = 6/9 and (1/3)(1/3) = 1/9 so x = 1/3

OpenStudy (anonymous):

9x^2 - 6x +1 = 0 (3x)^2 - 2*3x*1 + 1^2 = 0 (3x -1)^2 = 0